Given that X is inversely proportional to the square of Y. It is known that x = 8 for a particular value of Y. When Y is increased bu 200%, find
(a) the value of X,
(b) the percentage change in the value of X

Answer by Boreal(15235)

y^2=k/x. When x=8
y^2=k/8; k=8y^2
increasing y 200% is tripling
(3y)^2=9y^2, so y^2 is increased 9 fold
k=8y^2
9y^2=8y^2/x
x=8/9
The per cent change is the decrease over the original amount or 64/9 divided by 9 or 64/81 or 79%.